OpenStudy (anonymous):

Is anyone able to help me with a question regarding Mitosis in a human cell?

OpenStudy (anonymous):

Suppose you are working in a lab that grows various types of human body cells in petri dishes, a technique known as “cell culturing”. While you are looking at one particular cell culture under the microscope, you decide to count up the numbers of cells in the various phases of the mitotic cell cycle. You count 45 cells in anaphase, 34 cells in prophase, 23 cells in telophase, 11 cells in metaphase and 1008 cells in interphase. Could you use these numbers to determine the relative amounts of time this cell type spends in each phase of the mitotic cell cycle?

OpenStudy (aaronq):

yes, you could.

OpenStudy (anonymous):

would you be able to explain it to me excluding interphase because thats part of the cell cycle but not actual mitosis?

OpenStudy (aaronq):

They want the relative amount of time each phase assuming: 45+34+23+11+1008= 100% of the time i got those numbers from: "You count 45 cells in anaphase, 34 cells in prophase, 23 cells in telophase, 11 cells in metaphase and 1008 cells in interphase" Hm i don't know if they want you to include interphase. I mean it's part of the cell cycle, but it's definitely not mitosis. if you don't include interphase, then: 45+34+23+11= 100% of the time

OpenStudy (anonymous):

what do you mean 100% of the time?

OpenStudy (aaronq):

we'll they want a relative amount of time 45+34+23+11=113= 100% of the time (relative) so for example, time spent in anaphase 45/113*100%=39.8 % of the time This is if were not including interphase.

OpenStudy (anonymous):

oh okay so then the answer would be yes

OpenStudy (anonymous):

thank you so much!!!

OpenStudy (anonymous):

do you know the phases in order?

OpenStudy (anonymous):

nevermind! Thank you

OpenStudy (aaronq):

no problem!

OpenStudy (anonymous):

You forgot to give him a medal @junae54

OpenStudy (aaronq):

lol i don't care about medals, but thanks dude.

OpenStudy (anonymous):

i dont know how

OpenStudy (anonymous):

Click ''best response'' to aronqs answer.

OpenStudy (anonymous):

There we go!!!

OpenStudy (anonymous):

Awesome :3

OpenStudy (aaronq):

thanks !